# USED: Universal Speaker Extraction and Diarization

### v2

@@ Abstract
Speaker extraction and diarization are two enabling techniques for real-world speech applications. Speaker extraction aims to extract a target speaker’s voice from a speech mixture, while speaker diarization demarcates speech segments by speaker, annotating ‘who spoke when’. Previous studies have typically treated the two tasks independently. In practical applications, it is more meaningful to have knowledge about ‘who spoke what and when’, which is captured by the two tasks. The two tasks share a similar objective of disentangling speakers. Speaker extraction operates in the frequency domain, whereas diarization is in the temporal domain. It is logical to believe that speaker activities obtained from speaker diarization can benefit speaker extraction, while the extracted speech offers more accurate speaker activity detection than the speech mixture. In this paper, we propose a unified model called Universal Speaker Extraction and Diarization (USED) to address output inconsistency and scenario mismatch issues. It is designed to manage speech mixture with varying overlap ratios and variable number of speakers. We show that the USED model significantly outperforms the competitive baselines for speaker extraction and diarization tasks on LibriMix and SparseLibriMix datasets. We further validate the diarization performance on CALLHOME, a dataset based on real recordings, and experimental results indicate that our model surpasses recently proposed approaches.
@@ I Introduction
Each speaker has unique voice characteristics. Speaker extraction seeks to extract speech from the target speaker with reference to his/her voiceprint [1], while speaker diarization analyzes the audio signal to determine “who spoke when” [2]. As the important speech processing front-ends, they are widely used in various real-world speech applications, such as speaker verification [3, 4] and speech recognition [5, 6].
In a multi-talker scenario, the human brain processes incoming speech signals by performing speaker extraction and diarization simultaneously. Prior studies typically treated speaker extraction [7, 8, 9] and diarization [10, 11, 12, 13, 14] as two independent tasks. However, in practical applications, such as meeting analysis and mobile recording applications, solely relying on the results from one task is insufficient since it can only provide speaking boundaries (timestamp information) or clean speech (content information). Associating extracted speech and speaker labels offers more meaningful and interpretable results as it provides a comprehensive view of “who spoke what and when” [15]. Moreover, this association provides clean speech signals so that it can further facilitate the speaker-attributed automatic speech recognition task, which aims to answer “who spoke what” [16, 17, 18].
@@
Most of the previous studies dealt with speaker extraction and speaker diarization either independently [11, 12, 7, 9] or jointly under some controlled conditions [22, 23], such as with a fixed number of speakers or a pre-defined overlap ratio. To date, no single-model approach has been proposed to perform speaker extraction and speaker diarization for speech mixture with varying overlap ratios and variable number of speakers. This prompts us to look into effectively incorporating speaker extraction and speaker diarization to benefit from their interaction.
With this motivation, we propose the Universal Speaker Extraction and Diarization model (USED) to address output inconsistency and scenario mismatch. The term ‘universal’ is defined by two aspects: (a) the ability to perform for a variable number of speakers, and (b) the ability to process speech mixture with an arbitrary overlap ratio.
To this end, we design the model by considering two aspects: data ingress and data egress. From data ingress perspective, we design a novel embedding assignment module driven by speech references to mitigate output inconsistency naturally. This module yields two advantages. Firstly, it enables the generation of outputs for a variable number of speakers according to the speech references, thereby ensuring consistency across tasks. Secondly, it helps the model align speaker extraction and speaker diarization outputs to avoid output permutation problem [24]. From data egress perspective, we extend the scope of overlap ratio by a multi-task interaction module to alleviate the scenario mismatch between speaker extraction and speaker diarization. Specifically, the multi-task interaction module leverages speaker diarization results to mitigate background noise of the extracted speech and enhance the optimization of scenario-aware differentiated (SAD) loss [25] within our USED. Simultaneously, speaker extraction produces a cleaner speech signal to assist in the prediction of speaker diarization, particularly for the overlap part.
@@
  • We propose a universal solution, i.e. USED, that performs the speaker diarization and extraction jointly. The solution deals with speech mixture with an arbitrary number of speakers and diverse overlap ratios.
@@ II-A Speech Separation and Speaker Extraction
Recent studies on blind speech separation have made significant progress, with the development of advanced models such as TasNet [26], Dual-Path Recurrent Neural Network (DPRNN) [27], Band-Split Recurrent Neural Network (BSRNN) [28] and efficient training algorithms such as Permutation Invariant Training (PIT) [24]. However, speech separation methods suffer from the global permutation ambiguity problem [8], which makes it hard to process long utterances. Moreover, it requires prior knowledge or estimation of the number of speakers in the speech mixture in advance, which is hard to get in real-world applications. Unlike blind speech separation that seeks to separate all speakers, speaker extraction only extracts speech for one target speaker at a time. In general, speaker extraction methods are classified into frequency-domain methods, such as SpeakerBeam [29, 30, 31] and VoiceFilter [32], and time-domain methods, e.g. SpEx+ [7] and X-SepFormer [9]. Time-domain approaches can naturally avoid the phase estimation problem, which exists in frequency-domain approaches [7]. Therefore, we develop the USED model based on the time-domain approach, SpEx+.
It is common that speech separation and speaker extraction systems in the literature are optimized for highly overlapped speech mixture [26, 7, 9]. However, the overlap ratio in daily conversations and meetings is typically around 20% [20, 21], leading to a scenario mismatch between training on highly overlapped speech and evaluating on sparsely overlapped speech. To address such mismatch, speaker extraction for sparsely overlapped speech was recently studied [33, 34, 25]. It was suggested to minimize speech power when the target speaker is absent, while maximizing the signal-to-distortion ratio (SDR) or scale-invariant signal-to-distortion ratio (SI-SDR) [35] when the target speaker is present. This introduces another challenge that is hard to balance the two scenario-dependent objectives, i.e., SI-SDR and power, during training. We will study a multi-task interaction module in the USED model to overcome this challenge.
@@ II-B Speaker Diarization
Traditional clustering-based diarization methods [36, 37, 38, 39] typically comprise multiple components that are trained separately. These methods implicitly assume that each speech segment is only from one speaker, which fails when multiple people speak at the same time. Two mainstream studies are attempting to solve this problem, i.e. end-to-end neural diarization (EEND) [10, 11], and target speaker voice activity detection (TS-VAD) [12]. EEND seeks to minimize the diarization error directly with permutation-free objectives [10], while TS-VAD uses the speaker embedding from the clustering results as the reference to detect the speaking status of each person. These methods have obtained good performance and set the stage for our study.
@@ III-A Task Definition
Let $x$ be a multi-talker speech mixture with varying overlap ratio from 0% to 100 %, which consists of the speech signals from $l$ speakers and noise signal $n$. The speech mixture can be represented as,
@@
where $l>1$, and $c_{i}$ denotes the speech signal of speaker $i$. Each speaker $i$ has a corresponding speech reference, denoted as $z_{i}$, which is estimated from speech mixture or provided as pre-enrolled information.
Given a speech mixture and speech references, the goal of the USED model is to generate separated speech signals and speaker activity information concurrently for a variable number of speakers, ranging from $1$ to $l$. The number of speakers corresponds to the available speech references. This is motivated by scenarios where obtaining speech references for all speakers is challenging, or the target speakers are only a subset of all speakers. During inference, if there is only one target speaker, the USED model outputs a speech signal and diarization result for that speaker in a similar way to a speaker extraction model [7]. However, if there are $l$ target speakers, the model would output multiple speech signals and diarization results concurrently, each for one speaker, just like a speech separation model does [26].
@@ III-B System Overview
The speaker encoder converts the available speech references into speaker embeddings, that are the feature representation of the target speakers. Meanwhile, the speech encoder generates spectrum or spectrum-like feature representation from the speech mixture $x$.
Unlike TS-VAD [12], which generates diarization results for all speakers according to speaker embeddings and speech mixture, we propose an embedding assignment module to enable custom control over the number of model’s outputs. The embedding assignment module prepares the input embeddings with three different states, active, blank, and residual, based on the speaker embeddings, which are responsible for generating results accordingly. The active state is used as a position marker for the expected speakers. The blank state serves as a position marker to guide the network in outputting a silent segment. The residual state marks the residual unexpected speakers. The blank state is motivated by the Connectionist Temporal Classification (CTC) [41], which integrates blank characters to manage pauses in speech recognition and gaps between characters in Optical Character Recognition (OCR). The embedding with blank state is designed to clearly distinguish the outputs from the expected and unexpected speakers. The reason for this design is that the output from the unexpected speaker is usually a silent segment. With the help of the embedding assignment module, the USED model naturally avoids the output inconsistency stemming from the integration of speaker extraction and diarization.
@@
The USED model differs from typical speaker extraction and speaker diarization systems [10, 7] in two key aspects. Firstly, it outputs speech signals for an arbitrary number of speakers, ranging from $1$ to $l$, with the help of the embedding assignment module. Secondly, it seeks to effectively deal with speech mixture with a wide range of overlapping ratios.
Figure 1: The overall training and inference framework of our proposed USED model. It comprises a speech encoder, a speaker encoder, an embedding assignment module, a separator, an extraction decoder and a diarization decoder. The USED model generates both extracted speech and speaker diarzation results by leveraging the speech mixture and reference as input. The symbols $\otimes$ and $\ominus$ refer to element-wise multiplication and frame-wise concatenation, respectively. The TCN block and mask module are denoted by rounded rectangles, which are described in detail in Fig. 2 or introduced later. Different colours are used to distinguish network layers from different components. The active state, blank state and residual state are assigned by Algorithm 1.
@@ III-C Speech Encoder
Analogous to a frequency analyzer, the speech encoder aims to generate spectrum-like frame-based embedding sequences from speech mixture and speech reference. Inspired by SpEx+[7], our approach leverages a multi-scale speech encoder.
@@
where the function $e(\cdot)$ denotes the operation performed by the multi-scale speech encoder. The set $Z=\{Z_{1},...,Z_{l}\}$ represents the embedding sequences for $l$ speakers. The input and output channel sizes of three Conv1Ds are one and $C$, respectively. The kernel sizes of three Conv1Ds are denoted as $L^{e}_{1}$, $L^{e}_{2}$ and $L^{e}_{3}$, while they have a common stride of $\frac{L^{e}_{1}}{2}$. The outputs $X$ and $Z_{i}$ have a shape of $3\times T\times C$, where $T$ is the number of frames. For simplicity, Fig. 1 does not explicitly depict the multi-scale structure.
@@
Figure 2: The details of model structure for ResNet block and TCN layer. BN, PReLU, gLN and D-Conv are batch normalization, parametric ReLU, global layer normalization and dilated depth-wise separable convolution. The symbol $\oplus$ refers to the element-wise addition.
@@ III-D Speaker Encoder
Speaker encoder is designed to extract speaker embedding from the corresponding enrolled speech. Specifically, we employ a speaker encoder to extract speaker embeddings, denoted as $V_{1},...,V_{l}$, from the embedding sequences $Z$. These speaker embeddings capture the distinctive characteristics of each speaker. After layer normalization (LN), we leverage a CNN layer with a kernel size of $1\times 1$ ($1\times 1$ CNN) on the embedding sequence $Z_{i}$ for speaker $i$. This is subsequently followed by $N^{r}$ residual network (ResNet) blocks [42]. The input and output channel sizes of this $1\times 1$ CNN are $3C$ and $C$, respectively. Finally, an additional $1\times 1$ CNN, combined with a mean pooling operation, generates a speaker embedding $V_{i}$ for speaker $i$ with an embedding dimension $D_{\text{spk}}$. After that, the speaker encoder predicts speaker identities, denoted as $\hat{y}^{\text{spk}}_{i}$, through a linear layer followed by a softmax (SM) activation function for speaker $i$.
The detailed architecture of the ResNet we used is shown in Fig. 2a, a ResNet block comprises two $1\times 1$ CNNs and a max-pooling layer. Each $1\times 1$ CNN within the ResNet block is accompanied by batch normalization (BN) and parametric ReLU (PReLU) for normalization and non-linear transformation. A skip connection adds the input to the output obtained after the second BN layer. Lastly, the max-pooling layer is responsible for adjusting the length of the output embedding.

### v3

@@ Abstract
Speaker extraction and diarization are two enabling techniques for real-world speech applications. Speaker extraction aims to extract a target speaker’s voice from a speech mixture, while speaker diarization demarcates speech segments by speaker, annotating ‘who spoke when’. Previous studies have typically treated the two tasks independently. In practical applications, it is more meaningful to have knowledge about ‘who spoke what and when’, which is captured by the two tasks. The two tasks share a similar objective of disentangling speakers. Speaker extraction operates in the frequency domain, whereas diarization is in the temporal domain. It is logical to believe that speaker activities obtained from speaker diarization can benefit speaker extraction, while the extracted speech offers more accurate speaker activity detection than the speech mixture. In this paper, we propose a unified model called Universal Speaker Extraction and Diarization (USED) to address output inconsistency and scenario mismatch issues. It is designed to manage speech mixtures with varying overlap ratios and variable number of speakers. We show that the USED model significantly outperforms the competitive baselines for speaker extraction and diarization tasks on LibriMix and SparseLibriMix datasets. We further validate the diarization performance on CALLHOME, a dataset based on real recordings, and experimental results indicate that our model surpasses recently proposed approaches.
@@ I Introduction
Each speaker has unique voice characteristics. Speaker extraction seeks to extract speech from the target speaker with reference to his/her voiceprint [1], while speaker diarization analyzes the audio signal to determine ‘who spoke when’ [2]. As important speech processing front-ends, they are widely used in various real-world speech applications, such as speaker verification [3, 4] and speech recognition [5, 6, 7, 8].
In a multi-talker scenario, the human brain processes incoming speech signals by performing speaker extraction and diarization simultaneously. Prior studies typically treated speaker extraction [9, 10, 11] and diarization [12, 13, 14, 15, 16] as two independent tasks. However, in practical applications, such as meeting analysis and mobile recording applications, solely relying on the results from one task is insufficient since it can only provide speaking boundaries (timestamp information) or clean speech (content information). Associating extracted speech and speaker labels offers more meaningful and interpretable results as it provides a comprehensive view of ‘who spoke what and when’ [17]. Moreover, this association provides clean speech signals so that it can further facilitate the speaker-attributed automatic speech recognition task, which aims to answer ‘who spoke what’ [18, 19, 20].
@@
Most of the previous studies dealt with speaker extraction and speaker diarization either independently [13, 14, 9, 11] or jointly under some controlled conditions [24, 25], such as with a fixed number of speakers or a pre-defined overlap ratio. To date, no single-model approach has been proposed to perform speaker extraction and speaker diarization for speech mixtures with varying overlap ratios and variable number of speakers. This prompts us to look into effectively incorporating speaker extraction and speaker diarization to benefit from their interaction.
With this motivation, we propose the Universal Speaker Extraction and Diarization model (USED) to address output inconsistency and scenario mismatch. The term ‘universal’ is defined by two aspects: (a) the ability to perform for a variable number of speakers, and (b) the ability to process speech mixtures with an arbitrary overlap ratio.
To this end, we design the model by considering two aspects: data ingress and data egress. From a data ingress perspective, we design a novel embedding assignment module driven by speech references to mitigate output inconsistency naturally. This module yields two advantages. Firstly, it enables the generation of outputs for a variable number of speakers according to the speech references, thereby ensuring consistency across tasks. Secondly, it helps the model align speaker extraction and speaker diarization outputs to avoid the output permutation problem [26]. From a data egress perspective, we extend the scope of overlap ratio by a multi-task interaction module to alleviate the scenario mismatch between speaker extraction and speaker diarization. Specifically, the multi-task interaction module leverages speaker diarization results to mitigate background noise of the extracted speech and enhance the optimization of scenario-aware differentiated (SAD) loss [27] within our USED. Simultaneously, speaker extraction produces a cleaner speech signal to assist in the prediction of speaker diarization, particularly for the overlap parts.
@@
  • We propose a universal solution, i.e. USED, that performs the speaker diarization and extraction jointly. The solution deals with speech mixtures with an arbitrary number of speakers and diverse overlap ratios.
@@ II-A Speech Separation and Speaker Extraction
Recent studies on blind speech separation have made significant progress, with the development of advanced models such as Conv-TasNet [28], Dual-Path Recurrent Neural Network (DPRNN) [29], Band-Split Recurrent Neural Network (BSRNN) [30] and efficient training algorithms such as Permutation Invariant Training (PIT) [26]. However, speech separation methods suffer from the global permutation ambiguity problem [10], which makes it hard to process long utterances. Moreover, it requires prior knowledge or estimation of the number of speakers in the speech mixture in advance, which is hard to get in real-world applications. Unlike blind speech separation that seeks to separate all speakers, speaker extraction only extracts speech for one target speaker at a time. In general, speaker extraction methods are classified into frequency-domain methods, such as SpeakerBeam [31, 32, 33] and VoiceFilter [34], and time-domain methods, e.g. SpEx+ [9] and X-SepFormer [11]. Time-domain approaches can naturally avoid the phase estimation problem, which exists in frequency-domain approaches [9]. Therefore, we develop the USED model based on the time-domain approach, SpEx+.
It is common that speech separation and speaker extraction systems in the literature are optimized for highly overlapped speech mixtures [28, 9, 11]. However, the overlap ratio in daily conversations and meetings is typically around 20% [22, 23], leading to a scenario mismatch between training on highly overlapped speech and evaluating on sparsely overlapped speech. To address such mismatch, speaker extraction for sparsely overlapped speech was recently studied [35, 36, 27]. It was suggested to minimize speech power when the target speaker is absent, while maximizing the signal-to-distortion ratio (SDR) or scale-invariant signal-to-distortion ratio (SI-SDR) [37] when the target speaker is present. This introduces another challenge since it is hard to balance the two scenario-dependent objectives, i.e., SI-SDR and power, during training. We will study a multi-task interaction module in the USED model to overcome this challenge.
@@ II-B Speaker Diarization
Traditional clustering-based diarization methods [38, 39, 40, 41, 42] typically comprise multiple components that are trained separately. These methods implicitly assume that each speech segment is only from one speaker, which fails when multiple people speak at the same time. Two mainstream studies are attempting to solve this problem, i.e. end-to-end neural diarization (EEND) [12, 13], and target speaker voice activity detection (TS-VAD) [14]. EEND seeks to minimize the diarization error directly with permutation-free objectives [12], while TS-VAD uses the speaker embedding from the clustering results as the reference to detect the speaking status of each person. These methods have obtained good performance and set the stage for our study.
@@ II-C Interaction Between Speech Separation and Diarization
Two recent works [44, 45] have explored the joint training of speech separation and speaker diarization based on unsupervised speech separation approaches. PixIT [44] employs PIT loss to integrate the speaker diarization task with MixIT [46] and draws on the best-of-both-worlds framework [47, 48] to stitch different segments based on the speaker diarization results. On the other hand, Neural FCASA [45] combines speaker diarization with neural FCA [49, 50], eliminating the need for a separate speaker diarization step to mask source activities. Both approaches jointly address the two tasks under the premise of unsupervised speech separation, focusing on long-form audio and optimizing the final separation results in terms of diarization error rate (DER) and word error rate (WER). In comparison, our USED model emphasizes the performance of supervised speech separation and speaker diarization across different overlap ratios and scenarios.
@@ III-A Task Definition
Let $x$ be a multi-talker speech mixture with varying overlap ratio from 0% to 100 %, which consists of the speech signals from $l$ speakers and a noise signal $n$. The speech mixture can be represented as,
@@
where $l>1$, and $c_{i}$ denotes the speech signal of speaker $i$. Each speaker $i$ has a corresponding speech reference, denoted as $z_{i}$, which is estimated from the speech mixture or provided as pre-enrolled information. Note that no normalization or noise reduction techniques are applied to the raw speech signals.
Given a speech mixture and speech references, the goal of the USED model is to generate separated speech signals and speaker activity information concurrently for a variable number of speakers, ranging from $1$ to $l$. The number of speakers corresponds to the available speech references. This is motivated by scenarios where obtaining speech references for all speakers is challenging, or the target speakers are only a subset of all speakers. During inference, if there is only one target speaker, the USED model outputs a speech signal and diarization result for that speaker in a similar way to a speaker extraction model [9]. However, if there are $l$ target speakers, the model would output multiple speech signals and a diarization results concurrently, one for each speaker, just like a speech separation model does [28].
@@ III-B System Overview
The speaker encoder converts the available speech references into speaker embeddings, that are the feature representations of the target speakers. Meanwhile, the speech encoder generates spectrum or spectrum-like feature representations from the speech mixture $x$.
Unlike TS-VAD [14], which generates diarization results for all speakers according to speaker embeddings and the speech mixture, we propose an embedding assignment module to enable custom control over the number of model’s outputs. The embedding assignment module prepares the input embeddings with three different states, active, blank, and residual, based on the speaker embeddings, which are responsible for generating results accordingly. The active state is used as a position marker for the expected speakers. The blank state serves as a position marker to guide the network in outputting a silent segment. The residual state marks the residual unexpected speakers. The blank state is motivated by Connectionist Temporal Classification (CTC) [51], which integrates blank characters to manage pauses in speech recognition and gaps between characters in Optical Character Recognition (OCR). The embedding with blank state is designed to clearly distinguish the outputs from the expected and unexpected speakers. The reason for this design is that the output from an unexpected speaker is usually a silent segment. With the help of the embedding assignment module, the USED model naturally avoids the output inconsistency stemming from the integration of speaker extraction and diarization.
@@
The USED model differs from typical speaker extraction and speaker diarization systems [12, 9] in two key aspects. Firstly, it outputs speech signals for an arbitrary number of speakers, ranging from $1$ to $l$, with the help of the embedding assignment module. Secondly, it seeks to effectively deal with speech mixtures with a wide range of overlapping ratios.
Figure 1: The overall training and inference framework of our proposed USED model. It comprises a speech encoder, a speaker encoder, an embedding assignment module, a separator, an extraction decoder and a diarization decoder. The USED model generates both the extracted speech and speaker diarization results by leveraging the speech mixture and the speech references as input. The symbols $\otimes$ and $\ominus$ refer to element-wise multiplication and frame-wise concatenation, respectively. The TCN block and mask module are denoted by rounded rectangles, which are described in detail in Fig. 2 and introduced later. Different colours are used to distinguish network layers from different components. The active, blank and residual states are assigned by Algorithm 1.
@@ III-C Speech Encoder
Analogous to a frequency analyzer, the speech encoder aims to generate spectrum-like frame-based embedding sequences from the speech mixture and speech references. Inspired by SpEx+[9], our approach leverages a multi-scale speech encoder.
@@
where the function $e(\cdot)$ denotes the operation performed by the multi-scale speech encoder. The set $Z=\{Z_{1},...,Z_{l}\}$ represents the embedding sequences for $l$ speakers. The input and output channel sizes of the three Conv1Ds are one and $C$, respectively. The kernel sizes of the three Conv1Ds are denoted as $L^{e}_{1}$, $L^{e}_{2}$ and $L^{e}_{3}$, while they have a common stride of $\frac{L^{e}_{1}}{2}$. The outputs $X$ and $Z_{i}$ have a shape of $3\times T\times C$, where $T$ is the number of frames. For simplicity, Fig. 1 does not explicitly depict the multi-scale structure.
@@
Figure 2: Model structure for a ResNet block and a TCN layer. BN, PReLU, gLN and D-Conv are batch normalization, parametric ReLU, global layer normalization and dilated depth-wise separable convolution. The symbol $\oplus$ refers to element-wise addition.
@@ III-D Speaker Encoder
The speaker encoder is designed to extract speaker embeddings from the corresponding enrolled speech. Specifically, we employ a speaker encoder to extract speaker embeddings, denoted as $V_{1},...,V_{l}$, from the embedding sequences $Z$. These speaker embeddings capture the distinctive characteristics of each speaker. After layer normalization (LN), we leverage a CNN layer with a kernel size of $1\times 1$ ($1\times 1$ CNN) on the embedding sequence $Z_{i}$ for speaker $i$. This is subsequently followed by $N^{r}$ residual network (ResNet) blocks [52]. The input and output channel sizes of this $1\times 1$ CNN are $3C$ and $C$, respectively. Finally, an additional $1\times 1$ CNN, combined with a mean pooling operation, generates a speaker embedding $V_{i}$ for speaker $i$ with an embedding dimension $D_{\text{spk}}$. After that, the speaker encoder predicts speaker identities, denoted as $\hat{y}^{\text{spk}}_{i}$, through a linear layer followed by a softmax (SM) activation function for speaker $i$.
The detailed architecture of the ResNet we used is shown in Fig. 2a. A ResNet block comprises two $1\times 1$ CNNs and a max-pooling layer. Each $1\times 1$ CNN within the ResNet block is accompanied by batch normalization (BN) and parametric ReLU (PReLU) for normalization and non-linear transformation. A skip connection adds the input to the output obtained after the second BN layer. Lastly, the max-pooling layer is responsible for adjusting the length of the output embedding.
We choose ResNet [52] because it has become a prominent architecture in speaker verification systems [53, 54, 55, 56] and is widely adopted in speaker extraction models [9, 57]. Its key advantage lies in the residual connections, which link frame-level layers and mitigate the vanishing gradient problem, facilitating faster convergence during backpropagation [52]. These residual connections also enable deeper neural network construction, allowing ResNet-based models to outperform regular CNNs. As a result, using ResNet as the backbone for speaker encoders leads to more robust and higher-quality speaker embeddings, making it a natural choice for this task.